Envision that you have served as business manager of Media World for over 2 years. You have noticed that for the last 12 months the business has regularly had cash assets of $20,000 or more at the end of each month. You have found a 6-month certificate of deposit that pays 6% compounded monthly. To obtain this rate of interest, you must invest a minimum of $2,000. You have also found a high interest savings account that pays 3% compounded daily. Based on the cash position of the business at this time, assume that you decide to invest $4,000.

2. Assume that you decide to invest the $4,000 in the high-interest savings account.

a. What future value would you expect to receive at the end of 6 months? (14 points)

b. How much interest would the investment earn for the period? (14 points)

c. What would be the effective rate of the investment? (14 points)
1
Expert's answer
2020-04-20T10:25:27-0400

a)Calculate a more profitable option:

"FV=4000(1+\frac{0.06}{12})^6=4121.51" , for 6 month 6% compounded monthly


"FV=4000(1+\frac{0.03}{365})^{183}=4060.62" , for 6 month 3% compounded monthly


profitable first option

b) Interest earned = FV-PV=4121.51-4000=121.51

Interest earned = FV-PV=4060.62-4000=60.62

c)"Effective rate = \frac{i}{PV}\times\frac{12}{6}=\frac{121.51}{4000}\times\frac{12}{6}=0.060755" ,6.0755%